OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of four x minus three, over two.?

OpenStudy (zzr0ck3r):

\[f(x)=\frac{4x-3}{2}\\ \implies \\y=\frac{4x-3}{2}\] solve for x

OpenStudy (zzr0ck3r):

then switch the x's and y's with each other

OpenStudy (anonymous):

f−1(x) = The quantity of two x plus three, over four ?

OpenStudy (anonymous):

f−1(x) = two over the quantity of three minus four x

OpenStudy (zzr0ck3r):

I told you how to get the answer, if you have a question about that process then ask. I cant just give you answers. I dont need to do the hw. you do:)